OpenStudy (anonymous):

Identify this function as continuous or not continuous on (-infinity,infinity): y=lxl-4

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

I have the answer and it is continuous but im not sure how to get that...like how do you know it is continuous?

jimthompson5910 (jim_thompson5910):

well |x| = x if x >=0 or |x| = -x if x <0 so effectively, y = |x|-4 is made of the two equations y = x - 4 if x >= 0 y = -x-4 if x < 0 and they together form y = |x|-4

jimthompson5910 (jim_thompson5910):

since y = x-4 and y = -x - 4 are both continuous at any point on either line, this means that y = |x|-4 is continuous everywhere

OpenStudy (anonymous):

ohh that makes sense thank you!

jimthompson5910 (jim_thompson5910):

you're welcome